OpenStudy (anonymous):

Evaluate 80 × a-2 for a = 4.

OpenStudy (anonymous):

use correct parenthesis

OpenStudy (anonymous):

i can only assume 80a - 2

OpenStudy (praxer):

80*(4-2)= 80*2 =160

OpenStudy (praxer):

if it is of form 80(a-2)

OpenStudy (anonymous):

According to BODMAS rule we can solve this problem first 80*a then from that sub -2 80*4=320 then 320-2=318

OpenStudy (praxer):

if 80*a-2= 80*4-2= 320-2 =318

OpenStudy (kewlgeek555):

\[80 \times a-2\]Where a = 4, so the equation turns to...\[80 \times 4-2\]Then, you go on and remember to use PEMDAS:\[80 \times 4-2\]\[320-2\]\[318\]And you're done. Simple as that. ;]

OpenStudy (anonymous):

thanks

OpenStudy (kewlgeek555):

A pleasure working with you and thank-you for using OpenStudy's mathematics section. :}

OpenStudy (anonymous):

Don't forget to click on the button called "Best Answer" to give a medal to the person who helped you :)